kobajet:

Layna can paint a 350-square foot room in 4 hours. It takes Bebe 7 hours to paint the same room. Which equation can be used to find the time in hours, t, it takes Layna and Bebe to paint the room together?

Mercury:

answer choices (please post these next time) 1. 1/4t + 1/7t = 1 2. 1/4t * 1/7t = 1 3. 4t + 7t = 1 4. 4t - 7t = 1

Mercury:

for shared work problems, the general idea is to determine what fraction of the whole project each person does per hour. then multiply the fractions by t, to get the total amount of work they do in (t) hours. then add up the fractions, and set them equal to 1 (representing 1 whole project). with that in mind, which answer choice do you think is the best model of the situation?